### v2

# Put CASH on Bandits: A Max K-Armed Problem for Automated Machine Learning

Abstract
The Combined Algorithm Selection and Hyperparameter optimization (CASH) is a challenging resource allocation problem in the field of AutoML. We propose MaxUCB, a max $k$-armed bandit method to trade off exploring different model classes and conducting hyperparameter optimization. MaxUCB is specifically designed for the light-tailed and bounded reward distributions arising in this setting and, thus, provides an efficient alternative compared to classic max $k$-armed bandit methods assuming heavy-tailed reward distributions. We theoretically and empirically evaluate our method on four standard AutoML benchmarks demonstrating superior performance over prior approaches. We make our code and data available at https://github.com/amirbalef/CASH_with_Bandits.
1 Introduction
The performance of machine learning (ML) solutions is highly sensitive to the choice of algorithms and their hyperparameter configurations which can make finding an effective solution a challenging task. AutoML aims to reduce this complexity and make ML more accessible by automating these critical choices (Hutter et al., 2019; Baratchi et al., 2024). For example, Hyperparameter optimization (HPO) methods focus on finding well-performing hyperparameter settings given a resource constraint, such as an iteration count or a time limit. However, in practice, it is often unclear which ML model class would perform best on a given dataset (Bischl et al., 2025). The problem of jointly searching the model class and the appropriate hyperparameters has been coined CASH, Combined Algorithm Selection and Hyperparameter optimization (Thornton et al., 2013). As a prime example, on tabular data, a ubiquitous data modality (van Breugel and van der Schaar, 2024), the state-of-the-art ML landscape covers classic ML methods, ensembles of gradient-boosted decision trees and modern deep learning approaches (Kadra et al., 2021; Gorishniy et al., 2021, 2024; McElfresh et al., 2023; Kohli et al., 2024; Hollmann et al., 2023; Holzmüller et al., 2024).
Figure 1: (Left) MaxUCB (blue line) outperforms combined search (black line) to identify the best-performing model class (brown line). (Middle) The irregular distribution of the empirical performance of model classes is left-skewed, and a higher mean may not correspond to a higher maximum. (Right) MaxUCB selects for which model to run one iteration of HPO during two-level optimization.
A popular approach to address the CASH problem is to use categorical and conditional hyperparameters to run HPO directly on the combined hierarchical search space of models and hyperparameters. AutoML systems use this approach, which we call combined search, to search well-performing ML pipelines (Thornton et al., 2013; Feurer et al., 2015; Komer et al., 2014; Kotthoff et al., 2017; Feurer et al., 2022), but HPO remains inefficient in high-dimensional and hierarchical search spaces. A naive solution to address the scalability limitation is to run HPO independently for the smaller search spaces of each ML model class and then compare the found solutions. However, this solution often exceeds available computational resources and does not scale well with an increasing number of ML models. Figure 1 (left) illustrates the difference between searching each space individually (colored dashed lines) and combined search (black line) on an exemplary dataset.
To leverage the efficiency of HPO in low dimensions, we use a Multi-Armed Bandit (MAB) method (Lattimore and Szepesvári, 2020) to dynamically allocate our budget. As shown in Figure 1 (right), each time the bandit strategy pulls an arm, it runs one iteration of HPO to evaluate a new configuration, resulting in a loss. The negative of the loss is used as reward feedback for the bandit algorithm. This approach is known as two-level CASH (decomposed CASH)  (Hoffman et al., 2014; Liu et al., 2019).
While most classical MAB problems aim to maximize the average rewards over time, the goal of the bandit algorithm for two-level CASH should be to maximize the maximum reward observed over time: as illustrated in Figure 1 (middle), tuning the LightGBM (red) will eventually outperform CatBoost. This goal aligns with Max $K$-Armed Bandit (MKB) problems (Carpentier and Valko, 2014; Achab et al., 2017; Baudry et al., 2022), often referred to as Extreme Bandits.
We precisely address this open question through a thorough statistical analysis of the empirical reward distributions of HPO tasks. Our main contribution is a state-of-the-art algorithm for two-level CASH based on a novel extreme bandit algorithm we call MaxUCB (Algorithm 1). We demonstrate the performance of our method on four benchmarks, which highlights the relevance of our assumptions for a wide variety of CASH problems. We analyze the theoretical performance of MaxUCB (Theorem 4.2) through regret bounds that also justify our novel choice of exploration bonus for the type of distributions relevant to the CASH problem. Importantly, our objective is rather to propose a practical algorithm with good empirical performance on CASH rather than a general multi-purpose bandit algorithm, so our guarantees hold under carefully crafted assumptions that resolve previously open questions (Nishihara et al., 2016).
2 Solving CASH using Bandits
The CASH problem for supervised learning tasks is defined as follows (Thornton et al., 2013). Given a dataset ${\mathbb{D}}=\{D_{train},D_{valid}\}$ of a supervised learning task, let $\mathcal{A}=\{A^{(1)},...,A^{(K)}\}$ be the set of $K$ candidate ML algorithms, where each algorithm $A^{(i)}$ has its own hyperparameter search space $\boldsymbol{\Lambda}^{(i)}$. The goal is to search the joint algorithm and hyperparameter configuration space to find the optimal algorithm $A^{(i^{*})}$ and its optimal hyperparameter configuration $\boldsymbol{\lambda}^{*}$ that minimizes a loss metric $\mathcal{L}$, e.g., the validation error [^1] We note that $\mathcal{L}$ can also be the result of k-fold cross-validation or other evaluation protocols measuring the expected performance of a model on unseen data (Raschka, 2018).. Formally,
$A^{(i^{*})}_{\boldsymbol{\lambda}^{*}}\in\operatorname*{arg\,min}_{A^{(i)}\in\mathcal{A},\boldsymbol{\lambda}\in\boldsymbol{\Lambda}^{(i)}}\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}},{\mathbb{D}}).$ (1)
For our approach, we study the decomposed variant (Hoffman et al., 2014; Liu et al., 2019) and address the following two-level optimization problem depicted in Figure 1 (right): at the upper level, we aim to find the overall best-performing ML model $A^{(i^{*})}$ by selecting model $A^{(i)}\in\mathcal{A}$ iteratively, and at the lower level, we aim to find the best-performing configuration $\boldsymbol{\lambda^{*}}\in\boldsymbol{\Lambda}^{(i)}$ for the selected model $A^{(i)}$. Formally,
$A^{(i^{*})}\in\operatorname*{arg\,min}_{A^{(i)}\in\mathcal{A}}\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}^{*}},{\mathbb{D}}),\quad\text{s.t.}\quad\boldsymbol{\lambda}^{*}\in\operatorname*{arg\,min}_{\boldsymbol{\lambda}\in\boldsymbol{\Lambda^{(i)}}}\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}},{\mathbb{D}}).$ (2)
The right-hand side of Equation 2 in the lower level can be efficiently addressed by existing iterative HPO methods such as Bayesian optimization (BO) (Jones et al., 1998; Garnett, 2022), which has been demonstrated to perform well in practical settings (Snoek et al., 2012; Chen et al., 2018; Cowen-Rivers et al., 2022). BO fits a surrogate model and uses an acquisition function to find a promising configuration to evaluate next. On the upper level, or left-hand side of Equation 2, the challenge is to carefully allocate the budget $T$ of HPO runs to the $K$ models in a manner that trades off exploration of the hyperparameter space of all models and exploitation (optimization) of the most promising model. As already noted in previous work, this is a typical MAB problem (Cicirello and Smith, 2005; Streeter and Smith, 2006a; Nishihara et al., 2016; Metelli et al., 2022).
At time $t$, the bandit algorithm chooses model $I_{t}\in\mathcal{A}$, and we denote $\boldsymbol{\lambda}_{t}$ the configuration proposed by the HPO method in the lower level. As a reward $r_{i,t}$, we feed back to the bandit algorithm an evaluation of the negative loss:
$r_{i,t}=-\mathcal{L}(A^{(i)}_{\boldsymbol{\lambda}_{t}},{\mathbb{D}}).$ (3)
In general, and as opposed to standard MAB, this reward process is not i.i.d. conditionally on the arm choices because the loss of the models depends on the progress of HPO on each model class (arm) as well as additional loss evaluation noise. To be able to design a tractable bandit algorithm, it is crucial to find an appropriate way to model this process to build controllable estimators. We focus on this aspect in the next section.
To complete the bandit model of this problem, we need to choose a regret metric that defines the oracle objective we compare to, and indeed aligns with Equation 2. For HPO, the regret should target max-value objectives (Jamieson and Talwalkar, 2016; Nishihara et al., 2016).
$R(T)=\max_{k\leq K}\mathop{\mathbbm{E}[\max r_{k,t}]}_{t\leq T}-\mathop{\mathbbm{E}[\max r_{I_{t},t}]}_{t\leq T}.$ (4)
This regret describes the gap between the highest-possible oracle reward that could be obtained by pulling only the arm with the highest performance (left part) and the actual observed rewards obtained by applying our strategy (right part). Notably, the expectation is necessary to account for the inherent stochasticity of both the HPO procedure (e.g., random search or Bayesian optimization) and the ML models themselves (e.g., random initialization and training variability).
Instead of using MKB algorithms to directly address Equation 4, related prior works focus on alternative methods. For example, Hu et al. (2021) proposed and analyzed the Extreme-Region Upper Confidence Bounds (ER-UCB) algorithm maximizing the extreme region of the feedback distribution, assuming Gaussian rewards. More recently, Balef et al. (2024) have shown that existing MKB algorithms underperform when applied to two-level CASH, and they proposed methods for maximizing the quantile values instead of the maximum value.
As another alternative method, Li et al. (2020) framed the CASH problem as a Best Arm Identification (BAI) task and introduced the Rising Bandits algorithm (Li et al., 2020). This MAB method assumes that the reward function for each arm increases with each pull, following a rested bandit model with non-decreasing payoffs (Heidari et al., 2016) (which has been shown to have linear regret when the reward increment per pull exceeds a threshold (Metelli et al., 2022)). Rising Bandits can be used for our setting using the maximum observed performance of the HPO history as the reward. However, this algorithm assumes deterministic rewards and increasing concave reward functions. To weaken this assumption, Li et al. (2020) introduced a hyperparameter to increase initial exploration. Mussi et al. (2024) further weakened this assumption by assuming that the moving average of the rewards is an increasing concave function.
Generally, in BAI approaches, the goal is to identify the arm with the highest mean reward. Here, rewards are the result of HPO runs, so this objective does not align well with Equation 2: there is no reason to measure the quality of a model on average over a random subset of hyperparameters chosen by HPO. This approximation made by prior work is justified by the complexity of this modeling problem and the existence of solid foundations on BAI to build upon. But in this work, we propose a fully data-driven model that better fits the true CASH objective and results in better empirical results.
3 Data Analysis of HPO Tasks
The reward process in Equation 3 is complex as it depends on the model and on the chosen HPO algorithm. So, as discussed, it is necessary to model it. Rather than choosing a convenient parametric family, we conduct a thorough analysis of typical sequences of losses obtained on real benchmarks. For each ML model class (corresponding to arms in our setup), we run $T=200$ iterations of HPO, with $32$ repetitions (each using a different seed) on a varying number of datasets on four AutoML benchmarks (see Appendix D.2 and  C.1).
We first analyze the survival function of the reward distributions. Recall that for a random variable $X\sim d$, the survival function is defined by: $x\mapsto G(x)=P_{X\sim d}(X\geq x).$ Figure 2 shows the average empirical survival function of all observed performances (normalized between 0 and $1$ for each task) for each arm. We rank model classes (i.e., arms) based on their best performance per dataset and report results on two benchmarks (see Appendix C for more details and results on all benchmarks). We observe that the reward distributions are
  1. 1.
    bounded. The reward, which measures the performance of a model, is determined by a score metric, e.g., accuracy. The extreme values vary for each arm and depend on the capability of the model class and the complexity of the task. Therefore, even if we run HPO indefinitely, achieving an infinite reward is impossible. Consequently, each arm has a bounded support with different maximum values, and at least a single optimal arm exists. $\rightarrow$ We can define a sub-optimality gap (Definition  3.1).
  2. 2.
    short-tailed, left-skewed, and the right tail is not heavy. The rewards are concentrated near the maximal value per model class, and extreme events are not outliers. As the HPO method’s performance reaches a certain level, further optimization often yields only small gains as optima tend to have flat regions (Pushak and Hoos, 2022). Therefore, many configurations perform similarly well, resulting in a skewed distribution. [^2] This has also been observed for related tasks, e.g., neural architecture search on CIFAR-10 (Su et al., 2021). $\rightarrow$ We expect to observe many extreme rewards.
  3. 3.
    nearly stationary. This means that the optimal arm does not change over time. This is clear for TabRepoRaw. We observe significant changes in the distributions for YaHPOGym, but most of the sub-optimal arms remain sub-optimal over time. $\rightarrow$ Ranking of well-performing arms does not change over time.
YaHPOGym TabRepoRaw YaHPOGym TabRepoRaw
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: (Left) The average empirical survival function of the reward distribution per arm ranked per dataset. Thin lines correspond to segments of the reward sequence and show the distribution change over time. (Right) The average empirical survival function per dataset for the best and worst arm. Thin lines correspond to individual datasets.
Prior research has shown that existing MKB algorithms can underperform if their assumptions about the distributions are too weak or misaligned (Nishihara et al., 2016; Balef et al., 2024). This underperformance is largely due to our second observation, which significantly differs from the common assumptions of the existing algorithms.
Preliminaries. Based on these observations, we can formulate the following definition and assumption on which we develop and analyze our algorithm.
Definition 3.1.
The suboptimality gap $\Delta_{i}\geq 0$ for arm $i$ is defined as:
$\Delta_{i}=\mathbbm{E}\left[\max_{t\leq T}r_{i^{*},t}\right]-\mathbbm{E}\left[\max_{t\leq T}r_{i,t}\right]$
where $r_{i^{*},t}$ and $r_{i,t}$ are the rewards observed from optimal arm $i^{*}$ and arm $i$ at time $t$, respectively.
Assumption 3.2.
We assume that the i.i.d. random variable $X$, representing the rewards, follows a bounded distribution with support in $[a,b]$ and continuous survival function $G$.
Lemma 3.3.
Suppose Assumption 3.2 holds. Then, there exists $L,U\geq 0$ such that the survival function $G$ can be bounded near $b$ by two linear functions:
$\forall\epsilon\in(0,b-a),\quad L\epsilon\leq G(b-\epsilon)\leq U\epsilon.$ (5)
Lemma 3.3 (proof in Appendix B.1) provides a way to characterize the shape of any bounded distribution near its maximum value through distribution-dependent constants, $L$ and $U$ (see Appendix C.2 for more details and a visualization). Intuitively, it quantifies the behavior of the distribution of the ML model performance in a given hyperparameter space near the optimum. A large value of $L$ indicates a steep drop in the survival function near the maximum, while a small value of $U$ leads to a more gradual decay of the survival function and conversely (see Figure 3).
YaHPOGym TabRepoRaw
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 3: (Left, Middle) L and U for the average survival functions in Figure 2. (Right) We highlight the difference to right-skewed Log-normal and Pareto distributions.
YaHPOGym TabRepoRaw
[Uncaptioned image] [Uncaptioned image]
Figure 4: Histogram of $L$ and $U$ values across individual datasets.
We study the values of $L$ and $U$ in Lemma 3.3 on our empirical data and show the distribution of $L$ and $U$ in Figure 4 (see Appendix C.3 for more analyses). Next, we focus on the uniqueness of our assumptions before discussing how this impacts our MKB algorithm’s performance in Section 4.
Common Max $K$-armed Bandit assumptions are misaligned with CASH. Existing MKB algorithms can be classified into two main categories. First, distribution-free approaches do not leverage any assumption on the type of reward distributions (Streeter and Smith, 2006b; Bhatt et al., 2022; Baudry et al., 2022). Secondly, parametric and semi-parametric approaches typically assume that the reward distributions follow heavy-tailed distributions, typically following a second-order Pareto assumption whose parameters can be estimated using extreme value theory (Carpentier and Valko, 2014; Achab et al., 2017).
Our empirical analysis shows that existing MKB algorithms’ assumptions about reward distributions are either overly broad or misaligned with the CASH problem. We identify two key differences between our assumptions and those made by current MKB algorithms. First, the sub-optimality gap makes the regret definition (Equation 4) for bounded distributions meaningful. Without the existence of a nonzero gap, the regret definition fails since any policy consistently selecting a single arm can achieve zero regret, as Nishihara et al. (2016) pointed out with Bernoulli distributions. Second, our reward distributions differ from those used in existing MKB algorithms. Lemma 3.3 characterizes the shape of the distribution, with a higher $L$ ensuring more mass near extreme values, making extreme values easier to estimate. In our case, values for $L$ are mostly higher than $1$ while for heavy-tailed distributions, commonly used as the basis for MKB algorithms, are close to 0 (see the rightmost plot in Figure 3). In general, our analysis shows that considering the right range of $L$ values unlocks the problem raised by Nishihara et al. (2016), who focused on constructing counterexamples through the distributions that have unrealistic values for $L$. Thus, under our assumptions, their negative result [^3] Theorem $11$ in (Nishihara et al., 2016): “no policy can be guaranteed to perform asymptotically as well as an oracle that plays the single best arm for a given time horizon.”, which means any policy needs to explore all arms for budget T does not apply here.
4 MaxUCB
Based on Lemma 3.3 and the regret definition (Equation 4), we introduce MaxUCB in Algorithm 1 for $K$ arms with a limited time horizon of $T$ iterations. [^4] MaxUCB needs to store the number of pulls and the maximum reward for each arm, resulting in a memory requirement of $\mathcal{O}(K)$. The time complexity is $\mathcal{O}(KT)$.
Description of MaxUCB. Our algorithm balances exploration and exploitation according to the standard optimism principle at the heart of Upper Confidence Bound (UCB) bandit methods (Auer, 2002; Lattimore and Szepesvári, 2020). The main novelty we introduce is in the computation of a distribution-adapted exploration bonus for MaxUCB.
Our exploration bonus $(\frac{\alpha\log(t)}{n})^{2}$ deviates from typical UCB literature due to faster concentration of maximum values in bounded distributions. This is because the probability of bad events (violating confidence intervals for the expectation of the max, see Equation 30 in Appendix B.3) can be written as:
$P(\text{Bad events})\leq O\left(e^{-n\sqrt{C(n)}}+nC(n)\right),$ (6)
where distribution-dependent constants are hidden for clarity. Then, setting $C(n)=1/n^{2}$ minimizes the probability of bad events: the first term becomes independent of $n$, while the second term decreases with $n$. Notably, this faster concentration can only be obtained for the reasonably well-behaved distributions we consider following the study of the previous section and it is not a general property of the maxima; more details can be found in Appendix B.4. Furthermore, MaxUCB uses $\alpha\geq 0$ to control the exploration-exploitation trade-off; a higher $\alpha$ leads to more exploration.
Algorithm 1 MaxUCB
[…]
2:for each arm $i\leq K$ do
3:  Pull arm $i$, set $n_{i}\leftarrow 1$, observe reward $r_{i,1}$ $\triangleright$ Evaluating default configuration
4:end for
5:for $t=K+1$ to $T$ do
6:  for each arm $i\leq K$ do
7:    Update policy $U_{i}=\max{(r_{i,1},...,r_{i,n_{i}})}+(\frac{\alpha\log(t)}{n_{i}})^{2}$ $\triangleright$ differs from classical UCB, where $U_{i}=\bar{r}_{i}+\sqrt{\tfrac{\alpha\log(t)}{n_{i}}}$
8:  end for
9:  Select arm $I_{t}=\operatorname*{arg\,max}\limits_{i\leq K}U_{i}$ ,   $n_{I_{t}}\leftarrow n_{I_{t}}+1$ , then observe reward $r_{I_{t},n_{I_{t}}}$
10:end for
Analysis of MaxUCB. We first show a regret decomposition result specific to max $K$-armed bandits that directly relates the regret definition in equation 4 with the number of suboptimal trials.
Proposition 4.1.
(Regret Upper Bound) the regret upper bound up to time $T$ is related to the number of times sub-optimal arms are pulled:
$R(T)\leq\frac{\max\limits_{i\leq K}b_{i}}{T}\sum\limits_{i\neq i^{*}}^{K}N_{i}(T)$ (7)
where $N_{i}(T)=\mathbbm{E}(\sum\nolimits_{t=1}^{T}\mathbbm{1}{\{I_{t}=i\}})$ is the number of sub-optimal pulls of arm $i$, and $b_{i}$ is the upper bound on the support of the rewards of arm $i$.
The proof is provided in Appendix B.2 and relies on standard tools in the extreme bandit literature (Baudry et al., 2022). From this result, it is clear that we can now obtain an upper bound on the regret by controlling the number of suboptimal arm pulls $(N_{i}(T))_{i\neq i^{\star}}$ individually. Our main theoretical result below proves such an upper bound for Algorithm 1.
Theorem 4.2.
For any suboptimal arm $i\neq i^{\star}$, the number of suboptimal draws $N_{i}(T)$ performed by MaxUCB (Algorithm 1) up to time $T$ is bounded by
$N_{i}(T)\leq\frac{T^{1-2L_{i^{*}}\alpha\sqrt{\Delta_{i}}}}{1-2L_{i^{*}}\alpha\sqrt{\Delta_{i}}}+2\alpha\sqrt{U_{i}T}\log(T).$ (8)
The result of Theorem 4.2 (proof in Appendix B.3) highlights that MaxUCB primarily leverages two key properties: the sub-optimality gap $\Delta_{i}$ and the shape of the distribution, as defined in Lemma 3.3. Specifically, the performance improves with a larger sub-optimality gap $\Delta_{i}$ and higher values of $L_{i^{*}}$ ($L$ for the optimal arm), which means that samples drawn from the distribution of the optimal arm are likely to be close to the extreme values. Additionally, smaller values of $U_{i}$ ($U$ for a sub-optimal arm $i$), which means it is less likely to draw samples close to the extreme values, reduce the number selecting sub-optimal arm $i$, thus enhancing overall performance.
For our task, as is shown in Figure 3, the values for $L_{i^{*}}$ are higher than $1$ and $U_{i}$ less than $10$ in most cases, yielding high empirical performance. We compare the number of pulls observed in our experiments with the expected values based on our theoretical analysis in Appendix E.2, showing that our analysis is not loose. However, it is essential to note that, in general, finding the optimal arm is challenging if $\Delta_{i}$ is close to zero or $L_{i^{*}}$ is very small. We assess the performance of MaxUCB on synthetic tasks in Appendix E.4, showing that the performance of MaxUCB deteriorates on tasks that do not satisfy our assumptions.
Finally, we provide a regret upper bound that combines the decomposition in Proposition 4.1 with the individual upper bounds above. This requires finding a parameter $\alpha$ that resolves a trade-off between the arms and minimizes the total upper bound, as shown in Corollary 9, whose proof is immediate.
Corollary 4.3.
If $L_{i^{*}}$, $\min\limits_{i\neq i^{*}}(\Delta_{i})$, and $T$ are known in advance, then the total regret $R(T)$ can be bounded as follows by choosing the exploration parameter $\alpha$ appropriately:
$R(T)\leq\mathcal{O}\bigg(\frac{K\log T}{\sqrt{T}}\max\limits_{i\leq K}b_{i}\bigg),\quad\text{when }\alpha=\frac{1}{4L_{i^{*}}\sqrt{\min\limits_{i\neq i^{*}}\Delta_{i}}}\bigg(1-\frac{2\log(\log(T))}{\log(T)}\bigg).$ (9)
This final result helps to understand the role of $\alpha$ and serves as guidelines to choose it in practice. Specifically, Equation 9 shows that either a small value of $L_{i^{*}}$ or a small sub-optimality gap requires a higher value for $\alpha$. Intuitively, a small $L_{i}^{*}$ means that the max value of the best arm needs more samples to be nearly reached, and a small suboptimality gap means that the two best arms are close and so hard to distinguish. Indeed, these problem-dependent quantities are unknown to the practitioner, so a direct approach to calculate $\alpha$ is not feasible. Therefore, evaluating performance robustness under "loose tuning" of $\alpha$ is essential.
5 Performance on AutoML tasks
Finally, we examine the empirical performance of MaxUCB in an AutoML setting via reporting average ranking and the number of wins, ties, and losses across tasks for each benchmark (details in Appendix D.1). We first focus on the impact of the hyperparameter $\alpha$ and then compare our approach to others. Specifically, we show that our two-level approach performs better than single-level HPO on the joint space, and MaxUCB outperforms other state-of-the-art bandit methods. To begin, we will provide a brief overview of the experimental setup used across all experiments.
Experimental setup. We use four AutoML benchmarks [^5] We used the AutoML Toolkit (AMLTK) (Bergman et al., 2024). We ran HPO on a compute cluster equipped with Intel Xeon Gold 6 2406\,240 CPUs, requiring 20 00020\,000 CPU hours. We conducted the remaining experiments on a local machine with Intel Core i7-1370P, requiring an additional $32$ CPU hours. , all implementing CASH for tabular supervised learning differing in the considered ML models, HPO method, and datasets, which we detail in Table D.1. For TabRepo and Reshuffling, we use available pre-computed HPO trajectories, whereas for TabRepoRaw and YaHPOGym, we use SMAC (Hutter et al., 2011; Lindauer et al., 2022) as a Bayesian optimization method to run HPO ourselves.
Competitive Baselines. We compare against Quantile Bayes UCB (Balef et al., 2024), ER-UCB-S (Hu et al., 2021) and Rising Bandits (Li et al., 2020) which have been developed for the decomposed CASH task. We consider extreme bandits (QoMax-SDA (Baudry et al., 2022), Max-Median (Bhatt et al., 2022)) and classic UCB as general bandit methods. We use default hyperparameter settings for all methods. As combined search baselines, we consider Bayesian optimization (SMAC) and random search. [^6] Only available for TabRepoRaw and YaHPOGym, where we computed HPO trajectories ourselves. Additionally, we report the performance of the best (oracle) arm.
YaHPOGym TabRepoRaw
Refer to caption Refer to caption
Figure 5: The sensitivity of MaxUCB to hyperparameter $\alpha$, lower is better.
How sensitive is MaxUCB to the choice of $\alpha$? Figure 5, shows that the choice of $\alpha$ impacts performance. Lower values of $\alpha$ (light red lines) lead to good performance with small budgets, whereas higher values (dark red lines) achieve stronger final performance when sufficient time is available. An $\alpha$ around $0.5$ yields a balanced trade-off, offering robust anytime performance across tasks (see Appendix D.4 for detailed results). Another insight from this study is that the right choice of $\alpha$ may depend on characteristics of the datasets, such as the support of the losses, and could be meta-learned.
How does the two-level approach compare against combined search? We first compare the average rank over time in Figure 6 using combined search (black lines; random search and SMAC if available) to the two-level approach. We observe that all methods outperform random search (dotted line) and that while SMAC quickly catches up, some bandit methods continuously achieve a better (lower) rank. Additionally, we observe that most MAB algorithms (except Rising Bandits and QoMax-SDA) lead to superior performance in the early stages ($T=50$). This demonstrates that the decomposition is particularly useful when the number of iterations is limited. Additionally, Table 1 shows that the difference in final performance between combined search and the two-level approach is significant.
How does MaxUCB compare against other bandit methods? Figure 6 shows a substantial difference in the ranking. In the beginning, many methods perform competitively, but MaxUCB yields the best anytime and final performance. Classical UCB (red) and extreme bandits (QoMax-SDA; brown) perform worst. The Max-Median algorithm (purple) shows strong initial performance, but its effectiveness declines with more iterations. While Max-Median identifies and avoids the worst arm; it sometimes struggles to select the optimal arm, resulting in non-robust performance.
Next, we look at methods that were originally designed for AutoML. Both ER-UCB-S (pink) and Quantile Bayes UCB (orange) focus on estimating the higher region of the reward distribution rather than the extreme values. ER-UCB-S assumes a Gaussian reward distribution and is consistently outperformed by Quantile Bayes UCB, a distribution-free algorithm. Rising Bandits (green) underperforms initially due to its costly initialization but reaches a competitive final performance. This is especially pronounced for the YaHPOGym benchmark, where Rising Bandits outperforms MaxUCB with respect to normalized average performance (see Figure D.5 in Appendix D.5). This benchmark contains datasets where the optimal arm changes over time. Since Rising Bandits models non-stationary rewards, it performs better for these instances. [^7] A burn-in phase, i.e., pulling each arm for a few rounds at the beginning without observing the rewards, yields a competitive solution as we assess in Appendix E.3. Finally, MaxUCB and Quantile Bayes UCB are the only ones that significantly outperform combined search in Table 1. And looking at TabRepoRaw and Reshuffling, as depicted in Figure 6, demonstrates that MaxUCB is a robust method for CASH problems.
TabRepo[RS] TabRepoRaw[SMAC] YaHPOGym[SMAC] Reshuffling[HEBO]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Average rank of algorithms on different benchmarks, lower is better. SMAC and random search perform combined search across the joint space.
Benchmark MaxUCB Quantile Bayes UCB ER-UCB-S Rising Bandit QoMax-SDA Max-Median UCB
TabRepo [RS] p-value $\mathbf{\underline{0.00000}}$ $\mathbf{\underline{0.00000}}$ $\mathbf{\underline{0.00000}}$ $\mathbf{\underline{0.00000}}$ $\mathbf{\underline{0.00000}}$ $\mathbf{\underline{0.00006}}$ $\mathbf{\underline{0.00000}}$
w/t/l 186186/44/$10$ 179179/66/$15$ 135135/55/$60$ 185185/44/$11$ 172172/55/$23$ 126126/33/$71$ 135135/55/$60$
TabRepoRaw [SMAC] p-value $\mathbf{\underline{0.00072}}$ $\mathbf{\underline{0.00261}}$ $0.95063$ $0.42777$ $0.99194$ $0.99984$ $0.99194$
w/t/l 2424/0/66 2323/0/77 1111/0/1919 1616/0/$14$ 99/0/2121 66/0/$24$ 99/0/2121
YaHPOGym [SMAC] p-value $\mathbf{0.00880}$ $\mathbf{\underline{0.00503}}$ $0.31038$ $\mathbf{\underline{0.00074}}$ $0.08372$ $0.50000$ $\mathbf{0.02412}$
w/t/l 6464/0/$39$ 6565/0/$38$ 5454/11/$48$ 6868/0/$35$ 5959/0/$44$ 5252/0/$51$ 6262/0/$41$
Table 1: P-values from a sign test assessing whether bandit methods outperform combined search. P-values below $\alpha=0.05$ are underlined, while those below $\alpha=0.05$ after multiple comparison correction (adjusting $\alpha$ by #comparisons) are boldfaced, indicating that the two-level approach is superior to combined search. Additionally, we report the number of wins, ties, and losses (w/t/l).
6 Conclusions, Discussions and Future Work
This paper addresses the CASH problem, proposing MaxUCB, an MKB method. Our data-driven analysis answers an open question on the applicability of extreme bandits to CASH. We provide a novel theoretical analysis and show state-of-the-art performance on several important benchmarks. Limitations. Though our method can be applied beyond AutoML in principle, it is finely tuned for this setting with bounded and skewed distributions and for maximal value optimization. Our analysis relies on stationary distributions, which might not always be accurate, especially at the beginning of the HPO run, so a short burn-in phase may be needed to reach this regime. Our approach may not be distributionally optimal, as optimality in bounded extreme bandits remains an open question, and establishing lower bounds is left for future work. Lastly, we provide a default value for our hyperparameter $\alpha$ that might need adjusting for other applications.
Impact on AutoML systems. Our approach complements prior work on AutoML systems and increases their flexibility. First, our approach allows to choose any HPO method for each model at the lower level and thus it may integrate recent progress in HPO methods for some ML model classes, e.g., multi-fidelity or meta-learned methods  (György and Kocsis, 2011; Li et al., 2018; Falkner et al., 2018; Müller et al., 2023; Chen et al., 2022). Second, some AutoML systems (Swearingen et al., 2017; Li et al., 2023) decompose the search space into smaller subspaces to scale to a distributed setting and use Bandit methods to select promising subspaces (Levine et al., 2017; Li et al., 2020). While we focus on applying bandits to select promising ML models, our methods could also be applied in this setting. Finally, beyond CASH, MaxUCB is well-suited for sub-supernet selection in Neural Architecture Search (NAS) (Hu et al., 2022), showing similar reward distributions (Ly-Manson et al., 2024) (see Appendix E.5).
Choosing $\alpha$ adaptively. Figure 5 suggests that one could try to tune $\alpha$ online. However, it is known that, in theory, without additional information, data-adaptive parameters cannot be found at a reasonable exploration cost in bandit optimization settings (Locatelli and Carpentier, 2018). In AutoML systems, though, supplementary data, like estimates of the sub-optimality gap, reward distribution shape, and HPO convergence rate, can help adjust $\alpha$ adaptively.
Future Directions. To extend our extreme bandit setting, one could further refine the reward modeling by incorporating the non-stationarity, especially in AutoML for data streams, where optimal models shift with data distributions (Verma et al., 2024). Incorporating cost-aware optimization is another promising direction, as computational resources and time, rather than iteration counts, often define budgets in AutoML; this would require estimating model training times and factoring them into the decision process. Addressing the growing complexity of heterogeneous ML tasks, such as those involving pre-training, fine-tuning, or prompt engineering, may benefit from a hierarchical approach that allocates resources effectively across diverse (Balef and Eggensperger, 2025a, b). Additionally, exploiting structural similarities among algorithms and their hyperparameters could reduce exploration costs by sharing information across arms in the CASH problem, further enhancing efficiency.
Acknowledgments and Disclosure of Funding
The authors are funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC number 2064/1 – Project number 390727645. Additionally, C. Vernade acknowledges funding from the DFG under the project 468806714 of the Emmy Noether Programme. The authors also thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Appendix A Preliminiaries
Lemma A.1.
Let $X_{1},\ldots,X_{n}$ be $n$ samples independently drawn from distribution $d$, and let $G(x)=P(X>x)$ be the survival function. We have:
$P\left(\max_{1\leq t\leq n}X_{t}\leq x\right)$ $\leq e^{-nG(x)},$
$P\left(\max_{1\leq t\leq n}X_{t}>x\right)$ $\leq nG(x).$ (10)
Proof.
Let $F(x)=P(X\leq x)$ be the cumulative distribution function, so G(x)=1F(x)=P(X>x)G(x)=1-F(x)=P(X>x). First, consider the probability that the maximum of the $n$ samples is less than or equal to $x$:
$P\left(\max_{1\leq t\leq n}X_{t}\leq x\right)$ $=\prod_{i=1}^{n}P(X_{i}\leq x)=(F(x))^{n}=(1-G(x))^{n}\leq e^{-nG(x)},$ (11)
using the inequality $(1-x)^{n}\leq e^{-nx}$. Next, consider the probability that the maximum of the $n$ samples is greater than $x$:
$P\left(\max_{1\leq t\leq n}X_{t}>x\right)$ $\leq\sum_{i=1}^{n}P(X>x)=nG(x).$ (12)
Appendix B Proofs
B.1 Proof of Lemma 3.3
Lemma 3.3. Suppose Assumption 3.2 holds. Then, there exists $L,U\geq 0$ such that the survival function $G$ can be bounded near $b$ by two linear functions.
$\forall\epsilon\in(0,b-a),\quad L\epsilon\leq G(b-\epsilon)\leq U\epsilon$ (13)
Proof.
By applying the Mean Value Theorem (MVT) to the survival function $G$ over an interval $b-\epsilon,b]$, there exists a point $c\in(b-\epsilon,b)$ such that:
$G(b)-G(b-\epsilon)=G^{\prime}(c)(b-(b-\epsilon))=G^{\prime}(c)\epsilon$ (14)
Since $G^{\prime}(x)=-f(x)$ where $f(x)$ the probability density function (PDF) and $G(b)=0$ we have:
$G(b-\epsilon)=f(c)\epsilon$ (15)
Let $L$ and $U$ be the minimum and maximum values of the PDF $f(c)=\frac{G(b-\epsilon)}{\epsilon}$ over $\epsilon\in(0,b-a)$.
$L\epsilon\leq G(b-\epsilon)\leq U\epsilon$ (16)
Notably, in cases where we are interested in the survival function near some $b_{1}<b$ and $a_{1}>a$ i.e., over $(b_{1}-\epsilon,b_{1})$ applying the MVT again, there exists $c\in(b_{1}-\epsilon,b_{1})$ such that:
$G(b_{1})-G(b_{1}-\epsilon)=G^{\prime}(c)\epsilon=-f(c)\epsilon$ (17)
which rearranges to:
$G(b_{1}-\epsilon)=f(c)\epsilon+G(b_{1})$ (18)
For any $\epsilon\in(\delta,b_{1}-a_{1})$ with $\delta>0$ we can bound $G(b_{1}-\epsilon)$ as:
$L\epsilon\leq L\epsilon+G(b_{1})\leq G(b_{1}-\epsilon)\leq(f(c)+\frac{G(b_{1})}{\delta})\epsilon\leq U\epsilon$ (19)
B.2 Proof of Proposition 4.1
Proposition 4.1. (Upper Regret Bound) the upper regret bound up to time $T$ is related to the number of times sub-optimal arms are pulled,
$R(T)\leq\frac{\max\limits_{i\leq K}b_{i}}{T}\sum\limits_{i\neq i^{*}}^{K}N_{i}(T)$ (20)
Where $N_{i}(T)=\mathbbm{E}(\sum\limits_{t=1}^{T}\mathbbm{1}{\{I_{t}=i\}})$ is the number of sub-optimal pulls of arm $i$, arm $i^{*}$ is the optimal arm and $b_{i}$ is the upper bound on the reward of arm $i$ as given by Assumption 3.2, i.e., the reward of arm $i$ lies within the interval $a_{i},b_{i}]$.
Proof.
This proof is inspired by Assumption 1 of Baudry et al. [2022]. First, we need to determine an upper bound for the difference in the highest observed reward for arm $i$ when it has been pulled for $N_{i}(T)$ times compared to when it has been pulled for $T$ times.
$\mathbbm{E}\left[\max_{t\leq T}r_{i,t}\right]-\mathbbm{E}\left[\max_{t\leq N_{i}(T)}r_{i,t}\right]=\mathbbm{E}\left[\mathbbm{1}\left\{\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}=\max_{t\leq T}r_{i,t}\right\}\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}\right]$
$\leq\mathbbm{E}\Bigg[\mathbbm{1}\left\{\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}=\max_{t\leq T}r_{i,t}\right\}\underbrace{\mathbbm{1}\left\{\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}\leq B\right\}\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}}_{\leq B}\Bigg]$
$\quad+\mathbbm{E}\Bigg[\mathbbm{1}\left\{\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}=\max_{t\leq T}r_{i,t}\right\}\mathbbm{1}\left\{\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}>B\right\}\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}\Bigg]$
$\leq P\left(\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}=\max_{t\leq T}r_{i,t}\right)B$
$\quad+\mathbbm{E}\Bigg[\mathbbm{1}\left\{\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}=\max_{t\leq T}r_{i,t}\right\}\mathbbm{1}\left\{\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}>B\right\}\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}\Bigg].$ (21)
Since always $\max\limits_{N_{i}(T)+1\leq t\leq T}r_{i,t}\leq b_{i}$ by choosing $B=b_{i}$ we ensure that $\mathbbm{1}\left\{\max\limits_{N_{i}(T)+1\leq t\leq T}r_{i,t}>B\right\}=0$, leading to:
$\mathbbm{E}\left[\max_{t\leq T}r_{i,t}\right]-\mathbbm{E}\left[\max_{t\leq N_{i}(T)}r_{i,t}\right]$ $\leq P\left(\max_{N_{i}(T)+1\leq t\leq T}r_{i,t}=\max_{t\leq T}r_{i,t}\right)B$
$\leq\left(1-\frac{N_{i}(T)}{T}\right)B=\left(1-\frac{N_{i}(T)}{T}\right)b_{i}.$ (22)
Using this and according to the regret definition, we obtain the following:
$R(T)$ $=\mathop{\mathbbm{E}[\max r_{i^{*},t}]}_{t\leq T}-\mathop{\mathbbm{E}[\max r_{I_{t},t}]}_{t\leq T}$
$\leq\mathop{\mathbbm{E}\left[\max_{t\leq T}r_{i^{*},t}\right]}-\max_{i\leq K}\mathbbm{E}\left[\max_{t\leq N_{i}(T)}r_{i,t}\right]$
$=\min_{i\leq K}(\mathop{\mathbbm{E}\left[\max_{t\leq T}r_{i^{*},t}\right]}-\mathbbm{E}\left[\max_{t\leq N_{i}(T)}r_{i,t}\right])$
[…]
$\leq\min_{i\leq K}(\Delta_{i}+(1-\frac{N_{i}(T)}{T})b_{i})$
$\leq\min_{i\leq K}(\Delta_{i}+(1-\frac{N_{i}(T)}{T})\max\limits_{i\leq K}b_{i})$ (23)
Based on Definition 3.1, the suboptimality gap for the optimal arm $i^{*}$ is zero ($\Delta_{i^{*}}=0$ ). Additionally, the total number of pulls for the optimal arm $N_{i^{*}}(T)$ can be calculated as the difference between the total number of pulls across all arms $T$ and the pulls for all suboptimal arms ($i\neq i^{*}$), i.e. $N_{i^{*}}(T)=T-\sum\limits_{i\neq i^{*}}^{K}N_{i}(T)$. We upper bound the min in equation 23 by the specific value for the optimal arm $i^{*}$:
$R(T)\leq\frac{\max\limits_{i\leq K}b_{i}}{T}\sum\limits_{i\neq i^{*}}^{K}N_{i}(T)$ (24)
B.3 Proof of Theorem 4.2
Theorem 4.2. For any suboptimal arm $i\neq i^{\star}$, the number of suboptimal draws $N_{i}(T)$ performed by Algorithm 1 up to time $T$ is bounded by
$N_{i}(T)\leq\frac{T^{1-2L_{i^{*}}\alpha\sqrt{\Delta_{i}}}}{1-2L_{i^{*}}\alpha\sqrt{\Delta_{i}}}+2\alpha\sqrt{U_{i}T}\log(T)$ (25)
Proof.
In order to find the upper bound for the number of sub-optimal pulls of arm $i$ for the algorithm 1, without loss of generality, we assume that arm $1$ is the optimal arm, i.e. $i^{*}=1$. Let $\Delta_{i}=\mathbbm{E}[\max\limits_{t\leq T}r_{1,t}]-\mathop{\mathbbm{E}[\max\limits_{t\leq T}r_{i,t}]}$ be the suboptimality gap. Our goal is to determine an upper bound on $N_{i}(T)$, the number of times the sub-optimal arm $i$ has been pulled up to time $T$. First, we identify the event that the algorithm pulls the sub-optimal arm $i$ at time $t$:
$S$ $=\{\max{(r_{1,1},...,r_{1,n_{1}(t)})}+C_{t}(n_{1}(t))\leq\max{(r_{i,1},...,r_{i,n_{i}(t)})}+C_{t}(n_{i}(t))\}$
$=\{\max\limits_{1\leq n\leq n_{1}(t)}r_{1,n}+C_{t}(n_{1}(t))\leq\max\limits_{1\leq n\leq n_{i}(t)}r_{i,n}+C_{t}(n_{i}(t)\}$ (26)
where $S$ is the event of selecting a sub-optimal arm $i$ with a padding function $C_{t}(n)$. The exploration bonus $C_{t}(n)$ is a function that is designed to account for the exploration-exploitation trade-off and typically depends on $t$ and the number of times each arm has been pulled $n$. Let $n_{1}(t)$ and $n_{i}(t)$ represent the number of times the optimal arm $1$ and an sub-optimal arm $i$ have been pulled, respectively, where $n_{1}(t)\leq t$ and $n_{i}(t)\leq t$. We want to express $S$ in a union of events that covers all possible scenarios leading to $S$. Thus, we split $S$ into two complementary conditions as follows:
$S\subseteq\{\max\limits_{1\leq n\leq n_{1}(t)}r_{1,n}+C_{t}(n_{1}(t))\leq x\}\cup\{\max\limits_{1\leq n\leq n_{i}(t)}r_{i,n}+C_{t}(n_{i}(t))>x\}$ (27)
Where $x$ is a threshold value, we take $x=\mathbbm{E}[\max\limits_{t\leq T}r_{i,t}]$,
$S$ $\subseteq\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}+C_{t}(n_{1}(t))\leq\mathbbm{E}[\max_{t\leq T}r_{i,t}]\right\}$
$\quad\cup\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}+C_{t}(n_{i}(t))>\mathbbm{E}[\max_{t\leq T}r_{i,t}]\right\}$
$=\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-\Delta_{i}-C_{t}(n_{1}(t))\right\}$
$\quad\cup\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}-C_{t}(n_{i}(t))\right\}$ (28)
Thus, the event $S$ can be contained within the union of two bad events:
  • Underestimating the upper confidence bound of extreme values for the optimal arm 11
  • Overestimating the upper confidence bound of extreme values for the sub-optimal arm $i$
Now we use Lemma A.1 to calculate the probability of 28:
$P(S)$ $\leq P\left(\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathbbm{E}[\max_{t\leq T}r_{1,t}]-C_{t}(n_{1}(t))-\Delta_{i}\right\}\right)$
$\quad+P\left(\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathbbm{E}[\max_{t\leq T}r_{i,t}]-C_{t}(n_{i}(t))\right\}\right)$
$\leq e^{-n_{1}(t)G_{1}\left(\mathbbm{E}[\max\limits_{t\leq T}r_{1,t}]-C_{t}(n_{1}(t))-\Delta_{i}\right)}$
$\quad+n_{i}(t)G_{i}\left(\mathbbm{E}[\max_{t\leq T}r_{i,t}]-C_{t}(n_{i}(t))\right).$ (29)
Now, by applying Lemma 3.3, we can simplify the analysis by eliminating the complexities associated with survival functions $G_{1}$ and $G_{i}$.
$P(S)\leq e^{-n_{1}(t)L_{1}(C_{t}(n_{1}(t))+\Delta_{i})}+n_{i}(t)U_{i}C_{t}(n_{i}(t))$ (30)
For the first term of the right-hand side, by the Arithmetic Mean-Geometric Mean (AM-GM) inequality ($a+b\geq 2\sqrt{ab}$) of equation 30, we have:
$e^{-n_{1}(t)L_{1}(C_{t}(n_{1}(t))+\Delta_{i})}\leq e^{-2L_{1}n_{1}(t)\sqrt{C_{t}(n_{1}(t))\Delta_{i}}}$ (31)
In this stage, we want to find a proper padding function $C_{t}(n)$, which controls the right-hand side of equation 30 and equation 31. By choosing $C_{t}(n)=(\frac{\alpha\log(t)}{n})^{2}$, we have:
$e^{-n_{1}(t)L_{1}(C_{t}(n_{1}(t))+\Delta_{i})}$ $\leq e^{-2L_{1}n_{1}(t)\sqrt{C_{t}(n_{1}(t))\Delta_{i}}}=e^{-2L_{1}\alpha\log(t)\sqrt{\Delta_{i}}}=t^{-2L_{1}\alpha\sqrt{\Delta_{i}}}$ (32)
$n_{i}(t)U_{i}C_{t}(n_{i}(t))$ $\leq\frac{\alpha^{2}U_{i}\log^{2}(t)}{n_{i}(t)}$ (33)
This selection of the function of the exploration bonus results in two significant advantages. First, it provides an upper bound for the right-hand side of equation 30 that remains independent of $n_{1}$. Furthermore, Equation 31 shows a decreasing trend as the number of pulls for the sub-optimal arm $i$ increases. [^8] We note that this choice is not based on the inherent property of maximum values. In general one can use $C_{t}(n)=(\frac{\alpha\log(t)}{n})^{m}$ for $m>1$ with the optimal $m$ depending on the setting. In Appendix B.4 we show how $m$ affects the regret.
Now, we assume that the sub-optimal arm $i$ has been played for $l_{i}$ times, so $n_{i}(t)\geq l_{i}$. We want to calculate the number of sub-optimal pulls of arm $i$ up to time $T$:
$N_{i}(T)$ $\leq l_{i}+\sum_{t=l_{i}}^{T}P(S)\leq l_{i}+\sum_{t=l_{i}}^{T}t^{-2L_{1}\alpha\sqrt{\Delta_{i}}}+\sum_{t=l_{i}}^{T}\frac{\alpha^{2}U_{i}\log^{2}(t)}{l_{i}}$
$\leq l_{i}+\frac{T^{1-2L_{1}\alpha\sqrt{\Delta_{i}}}}{1-2L_{1}\alpha\sqrt{\Delta_{i}}}+\frac{\alpha^{2}U_{i}}{l_{i}}T\log^{2}(T)$ (34)
By choosing $l_{i}=\alpha\sqrt{U_{i}T}\log(T)$, we have:
$N_{i}(T)$ $\leq\frac{T^{1-2L_{1}\alpha\sqrt{\Delta_{i}}}}{1-2L_{1}\alpha\sqrt{\Delta_{i}}}+2\alpha\sqrt{U_{i}T}\log(T)$ (35)
B.4 Extension of Proof of Theorem 4.2
As we discuss in Section 3, the reward distribution in our setting is left-skewed. We now show that under the assumption that the survival function decays rapidly near the maximum i.e., $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])=\mathcal{O}(\frac{1}{T^{2}})$, a tighter bound can be derived. This assumption means that the reward distribution has a very light-tail near its upper extreme, which may often hold for our left-skewed distributions. Furthermore, we generalize our algorithm by assuming $C_{t}=(\frac{\alpha\log(t)}{n_{i}})^{m}$ as a exploration bonus function, where $m\geq 1$ is a hyperparameter.
Proof.
We begin with Equation 27 and we take $x=\mathbbm{E}[\max\limits_{t\leq T}r_{i,t}]+c\Delta_{i}$, where $c$ is an arbitrary variable $c\in[0,1]$. We have:
$S$ $\subseteq\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}+C_{t}(n_{1}(t))\leq\mathbbm{E}[\max_{t\leq T}r_{i,t}]+c\Delta_{i}\right\}$
$\cup\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}+C_{t}(n_{i}(t))>\mathbbm{E}[\max_{t\leq T}r_{i,t}]+c\Delta_{i}\right\}$
$=\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-(1-c)\Delta_{i}-C_{t}(n_{1}(t))\right\}$
$\cup\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}-C_{t}(n_{i}(t))+c\Delta_{i}\right\}$
$=\underset{S_{1}}{\underbrace{\left\{\max_{1\leq n\leq n_{1}(t)}r_{1,n}\leq\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-(1-c)\Delta_{i}-C_{t}(n_{1}(t))\right\}}}$ (36)
$\cup\underset{S_{2}}{\underbrace{\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}-C_{t}(n_{i}(t))+c\Delta_{i},\quad C_{t}(n_{i}(t))\leq c\Delta_{i}\right\}}}$ (37)
$\cup\underset{S_{3}}{\underbrace{\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}-C_{t}(n_{i}(t))+c\Delta_{i},\quad C_{t}(n_{i}(t))>c\Delta_{i}\right\}}}$ (38)
First, we calculate the probability of event $S_{2}$:
$P(S_{2})\leq P(\left\{\max_{1\leq n\leq n_{i}(t)}r_{i,n}>\mathop{\mathbbm{E}[\max_{t\leq T}r_{i,t}]}\right\})\leq n_{i}G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])\leq TG_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}).$ (39)
By calculating the number of sub-optimal pulls of arm $i$ up to time $T$, we know the third part of the event ($S_{3}$) can happen at most $C_{T}^{-1}(c\Delta_{i})=(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}$ times:
$N_{i}(T)$ $\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+\sum_{t=1}^{T}P(S_{1})+\sum_{t=1}^{T}P(S_{2})$ (40)
$\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+T^{2}G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])+\sum_{t=1}^{T}P(S_{1})$ (41)
$\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+M+\sum_{t=1}^{T}P(S_{1})$ (42)
Where $M$ is a constant as we assume $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])=\mathcal{O}(\frac{1}{T^{2}})$. Finally, we need to find an upper bound for $P(S_{1})$. We need to differentiate between two situations, when $m=1$ and when $m>1$
For $m=1$. We set $c=1$ and then we have:
$P(S_{1})$ $\leq e^{-n_{1}G_{1}(\mathop{\mathbbm{E}[\max_{t\leq T}r_{1,t}]}-(1-c)\Delta_{i}-C_{t}(n_{1}(t)))}\leq$ (43)
$e^{-L_{1}(C_{t}(n_{1}(t))+(1-c)\Delta_{i})}\leq e^{-\alpha L_{1}\log(T)}\leq T^{-\alpha L_{1}}$ (44)
Finally, we have:
$N_{i}(T)$ $\leq M+\frac{\alpha\log(T)}{\Delta_{i}}+\frac{T^{1-\alpha L_{1}}}{1-\alpha L_{1}}$ (45)
With $\alpha>\frac{1}{L_{1}}$:
$N_{i}(T)=\mathcal{O}(\frac{\log(T)}{L_{1}\Delta_{i}})$ (46)
For $m>1$. We know $n((\frac{a}{n})^{m}+b)\geq ab^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]$. We have:
$P(S_{1})$ $\leq e^{-n_{1}(t)L_{1}(C_{t}(n_{1}(t))+(1-c)\Delta_{i})}\leq e^{-\alpha L_{1}\log(T)((1-c)\Delta_{i})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}$ (47)
$\leq T^{-\alpha L_{1}((1-c)\Delta_{i})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}$ (48)
For simplicity, we set $c=\frac{1}{2}$. Then:
$N_{i}(T)$ $\leq(\frac{\alpha\log(T)}{c\Delta_{i}})^{\frac{1}{m}}+M+\sum_{t=1}^{T}P(S_{2})$ (49)
$\leq(\frac{2\alpha\log(T)}{\Delta_{i}})^{\frac{1}{m}}+M+\frac{T^{1-\alpha L_{1}(\frac{\Delta_{i}}{2})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}}{1-\alpha L_{1}(\frac{\Delta_{i}}{2})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}$ (50)
Finally, we have:
$N_{i}(T)=(\frac{2\alpha\log(T)}{\Delta_{i}})^{\frac{1}{m}}+\mathcal{O}(T^{1-\alpha L_{1}(\frac{\Delta_{i}}{2})^{\frac{m-1}{m}}[m(m-1)^{\frac{1}{m}-1}]}).$ (51)
And by choosing $\alpha$ as below
$\alpha=\mathcal{O}\left(\frac{1}{L_{1}(\Delta_{i})^{\frac{m-1}{m}}}\right),$ (52)
We have:
$N_{i}(T)=\mathcal{O}\left(\frac{\log(T)}{L_{1}\Delta_{i}^{{\frac{2m-1}{m}}}}\right)^{\frac{1}{m}}.$ (53)
We would like to emphasize again that this result only holds when $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])$ is sufficiently small. By using a weaker assumption, controlling $G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])$ is necessary. In Theorem B.3, we address this for our method by considering $C_{t}(n_{i}(t))\geq G_{i}(\mathbbm{E}[\max_{t\leq T}r_{i,t}])$ to ensure proper control.
Notably, in general, $\Delta_{i}$ depends on $T$ (see Definition 3.1). Meaning that $\frac{\log(T)}{\Delta_{i}}$ does not necessarily lead to a logarithmic regret. However, in some special scenarios where the reward distributions have different supports, we can ensure that $\Delta_{i}=b_{1}-b_{i}$ as $T$ approaches infinity, and a logarithmic regret is achievable.
Furthermore, Equation 53 shows that the parameter $m$ controls $N_{i}(T)$, the number of times a suboptimal arm is pulled asymptotically. When $T$ is very large, a higher value of $m$ (along with the optimal $\alpha$) improves performance. However, it makes the algorithm more sensitive to the choice of $\alpha$. As shown in Equation 52, with increasing $m$, $\Delta_{i}$ has a greater influence on the optimal $\alpha$. Noting that $\Delta_{i}$ varies across arms and is typically unknown for an unseen task, finding the optimal $\alpha$ is not feasible in practice. Therefore, there is a trade-off between performance and sensitivity. In the CASH setting, $m=2$ performs empirically well, while exhibiting low sensitivity to $\alpha$. Furthermore, as shown in Appendix D.4, we can find a range of $\alpha$ for which the MaxUCB works well across different CASH tasks.
Appendix C More Details on Reward Distribution
C.1 Reward distribution analysis
In addition to the analysis in the main paper in Figure 2 in Section 3, we collected the observed rewards (the output of HPO) for all arms on each model class. We calculate each dataset’s empirical survival function $G$ and provide the reward distribution analysis for all benchmark tasks. Notably, the shift in distribution (indicated by the thin lines) is low for all tasks, not contradicting the i.i.d. assumptions. For our method, we design our algorithm based on analyzing the distribution of raw rewards (in contrast to the distribution of maximum values over time).
Over time, the maximum value of samples generated from an i.i.d. distribution has an increasing trend, i.e., the extreme values get better. Notably, this is not contradictory with the Rising Bandits strategy Liu et al. [2019], which assumes that the maximum observed value over time is not decreasing (and then analyses the trend of this maximum observed value as a (non i.i.d) reward).
TabRepo
Refer to caption Refer to caption Refer to caption
TabRepoRaw
Refer to caption Refer to caption Refer to caption
YaHPOGym
Refer to caption Refer to caption Refer to caption
Reshuffling
Refer to caption Refer to caption Refer to caption
Figure C.1: (Left) The average empirical survival function of rewards (observed performances) per arm ranked per dataset. We divided the reward sequence into five segments over the budget (time horizon) to show the distribution change over time. Thin lines correspond to empirical survival functions for different segments, visualizing the change over time. (Middle) The average empirical survival function per dataset for the best and worst arm with thin lines corresponding to individual datasets. (Right) The sub-optimality gap $\Delta_{i}$.
C.2 More Details on Lemma 3.3
$L$ and $U$ are lower and upper bounds for the tangent line approximation of the survival function $G$ near the maximum value, indicating the shape of the distribution. We provide three examples to demonstrate this numerically.
Toy example: Assume two simple survival functions $G_{1}(x)=1-x^{2}$ (left skewed, blue curve) and $G_{2}(x)=(1-x)^{2}$ (right skewed, orange curve) with support $[0,1]$. We calculate $G(1-\epsilon)/\epsilon$ over some values of $\epsilon$ in the Figure C.2. To compute L and U, we determine the minimum and maximum values of $G(1-\epsilon)/\epsilon$ over the range $0<\epsilon<1$. For clarity, we restricted $\epsilon$ to iterate only over the set $\{0.1,0.3,0.5,0.7,0.9\}$. It implies that the calculated values of $L$ and $U$ are valid for $\epsilon\in[0.1,0.9]$.
$\epsilon$ $\frac{G_{1}(1-\epsilon)}{\epsilon}$ $\frac{G_{2}(1-\epsilon)}{\epsilon}$
0.10 1.90 0.10
0.30 1.70 0.30
0.50 1.50 0.50
0.70 1.30 0.70
0.90 1.10 0.90
L 1.10 0.10
U 1.90 0.90
Figure C.2: (Left) Determining $L$ and $U$ for $G_{1}$ and $G_{2}$. We calculate $\frac{G(1-\epsilon)}{\epsilon}$ over different ranges of $\epsilon$. The minimum and maximum values obtained from this ratio are assigned as $L$ and $U$, respectively. (Right) Showing two survival functions $G_{1}$ and $G_{2}$ along with their linear line approximations (gray lines). These tangent lines illustrate how $L$ and $U$ effectively bound the survival function $G$ near its maximum value.
Truncated uniform distribution: Assume that we have a truncated uniform distribution with support $[a,b]$. We know $G(x)=\frac{b-x}{b-a}$ for $x\in(a,b)$. For every $\epsilon\in[a,b]$ we have $\frac{G(b-\epsilon)}{\epsilon}=\frac{1}{b-a}$, which means $L=U=\frac{1}{b-a}$.
Truncated Gaussian distribution: There is no closed-form solution to formulate $L$ and $U$ based on the parameters of the truncated Gaussian distribution. Thus, we show the results of simulations to estimate $L$ and $U$ for truncated Gaussian within [0, 1] with various values $\mu$ and $\sigma$, averaging over $1000$ runs in Figure C.3.
$\mu$ $\sigma$ $L$ $U$
0.25 0.5 $0.58\pm 0.06$ $1.70\pm 0.48$
0.50 0.5 $0.85\pm 0.07$ $1.53\pm 0.34$
0.75 0.5 $1.01\pm 0.01$ $1.64\pm 0.25$
0.25 0.2 $0.34\pm 0.07$ $1.54\pm 0.28$
0.50 0.2 $0.44\pm 0.08$ $1.36\pm 0.04$
0.75 0.2 $1.01\pm 0.00$ $1.95\pm 0.04$
Figure C.3: (Left) Determining L and U for truncated Gaussian within [0, 1] with different values for $\mu$ and $\sigma$. Averaged over 1 0001\,000 runs. (Right) Showing survival function of truncated Gaussian distribution with different values for $\mu$ and $\sigma$.
C.3 Empirical Validation of Lemma 3.3
To study the empirical values for $L$ and $U$ in Lemma 3.3, we leverage the calculated empirical survival function $G$ for each dataset in Appendix C.1. Specifically, we evaluate $\frac{G(b-\epsilon)}{\epsilon}$ over the range $G^{-1}(0.99)<\epsilon<G^{-1}(0.01)$, where $G^{-1}(x)$ denotes the inverse of the survival function $G(x)$. Focusing on this range allows us to achieve a more robust estimation. Additionally, for TabRepo dataset, we exclude 23 datasets containing an arm with a standard deviation smaller than $0.001$, further enhancing the robustness of our analysis. In Figures C.4, C.5, C.6, and C.7, the evaluated values of $\frac{G(b-\epsilon)}{\epsilon}$ for different benchmarks are shown. Additionally, the values for $L$ and $U$, corresponding to $\min\limits_{\epsilon}(\frac{G(b-\epsilon)}{\epsilon})$ and $\max\limits_{\epsilon}(\frac{G(b-\epsilon)}{\epsilon})$, respectively, are presented. Finally, the histograms of these two variables are also included.
TabRepo
Figure C.4: Arms are ordered by sub-optimality gap. (Top) Thin orange lines represent $\frac{G(b-\epsilon)}{\epsilon}$, while the blue and red points correspond to $L$ and $U$ for our empirical reward distributions (see Lemma 3.3 for details). (Middle) Histogram of values for $L$. (Bottom) Histogram of values for $U$.
TabRepoRaw
Figure C.5: Arms are ordered by sub-optimality gap. (Top) Thin orange lines represent $\frac{G(b-\epsilon)}{\epsilon}$, while the blue and red points correspond to $L$ and $U$ for our empirical reward distributions (see Lemma 3.3 for details). (Middle) Histogram of values for $L$. (Bottom) Histogram of values for $U$.
YaHPOGym
Figure C.6: Arms are ordered by sub-optimality gap. (Top) Thin orange lines represent $\frac{G(b-\epsilon)}{\epsilon}$, while the blue and red points correspond to $L$ and $U$ for our empirical reward distributions (see Lemma 3.3 for details). (Middle) Histogram of values for $L$. (Bottom) Histogram of values for $U$.
Reshuffling
Figure C.7: Arms are ordered by sub-optimality gap. (Top) Thin orange lines represent $\frac{G(b-\epsilon)}{\epsilon}$, while the blue and red points correspond to $L$ and $U$ for our empirical reward distributions (see Lemma 3.3 for details). (Middle) Histogram of values for $L$. (Bottom) Histogram of values for $U$.
Appendix D More Details on the Experiments
D.1 Metric calculation
Average ranking calculation. We use bootstrapping with Monte Carlo sampling to calculate the average ranking plot with confidence intervals. For each time step and each task in every dataset, we resample the performance of each repetition (with replacement) and compute the average performance. We then rank the algorithms based on these averaged performances and repeat this process for all tasks. Finally, we average the rankings across tasks. This entire procedure was repeated 1000 times to estimate the confidence interval.
Number of wins, ties, and losses. To determine the number of wins, ties, and losses for each task in every dataset, we first compute the average performance of each algorithm over all repetitions at the final time step. We then perform pairwise comparisons of these averaged performances among all algorithms versus combined search. To account for negligible differences that are not statistically significant, we consider two performances to be tied if they are sufficiently close. Specifically, we use NumPy’s isclose function to compare the averaged performances, treating values within a default tolerance of $1\times 10^{-8}$ as equal.
D.2 Experimental Setup
Here, we provide details on our experimental setups as described in Table D.1.
name #models #tasks type HPO meth. (rep.) budget reference
YaHPOGym 6 103 surrogate SMAC ($32$) 200 [Pfisterer et al., 2022]
TabRepo 7 200 tabular random search ($32$) 200 [Salinas and Erickson, 2024]
TabRepoRaw 7 30 raw SMAC ($32$) 200 -
Reshuffling 4 10 raw HEBO ($30$) 250 [Nagler et al., 2024]
Table D.1: Overview of AutoML tasks. For TabRepo and Reshuffling, we use pre-computed HPO trajectories. TabRepoRaw resembles the same model space as TabRepo, but instead of random search, we run HPO ourselves. Similarly, we run HPO across provided surrogate HPO benchmark tasks YaHPOGym. We use SMAC [Lindauer et al., 2022, Hutter et al., 2011] implementing Bayesian optimization using Random forests for both tasks.
YaHPOGym [Pfisterer et al., 2022], a surrogate benchmark, covers $6$ ML models (details in Table D.2) on $103$ datasets and uses a regression model (surrogate model) to predict performances for queried hyperparameter settings. We use Bayesian optimization as implemented by SMAC [Lindauer et al., 2022] using Random Forests to conduct HPO. Additionally, we compare our two-level approach to combined search using SMAC, SMAC without initial design (SMAC-no-init), and Random Search. We run $32$ repetitions and use a budget of $200$ iterations for each evaluation.
Table D.2: Hyperparameter spaces for ML models in YaHPOGym.
ML model Hyperparameter Type Range Info
- trainsize continuous [0.03, 1] =0.525 (fixed)
imputation categorical {mean, median, hist} =mean (fixed)
Glmnet alpha continuous [0, 1]
s continuous [0.001, 1097] log
Rpart cp continuous [0.001, 1] log
maxdepth integer [1, 30]
minbucket integer [1, 100]
minsplit integer [1, 100]
SVM kernel categorical {linear, polynomial, radial}
cost continuous [4.5e-05, 2.2e4] log
gamma continuous [4.5e-05, 2.2e4] log, kernel
tolerance continuous [4.5e-05, 2] log
degree integer [2, 5] kernel
AKNN k integer [1, 50]
distance categorical {l2, cosine, ip}
M integer [18, 50]
ef integer [7, 403] log
ef_construction integer [7, 403] log
Ranger num.trees integer [1, 2000]
sample.fraction continuous [0.1, 1]
mtry.power integer [0, 1]
respect.unordered.factors categorical {ignore, order, partition}
min.node.size integer [1, 100]
splitrule categorical {gini, extratrees}
num.random.splits integer [1, 100] splitrule
XGBoost booster categorical {gblinear, gbtree, dart}
nrounds integer [7, 2980] log
eta continuous [0.001, 1] log, booster
gamma continuous [4.5e-05, 7.4] log, booster
lambda continuous [0.001, 1097] log
alpha continuous [0.001, 1097] log
subsample continuous [0.1, 1]
max_depth integer [1, 15] booster
min_child_weight continuous [2.72, 148.4] log, booster
colsample_bytree continuous [0.01, 1] booster
colsample_bylevel continuous [0.01, 1] booster
rate_drop continuous [0, 1] booster
skip_drop continuous [0, 1] booster
TabRepo [Salinas and Erickson, 2024] consists of pre-evaluated performance scores for $200$ iterations of random search for $7$ ML models (details in Table D.3) on $200$ datasets (context name: D244_F3_C1530_200D244\_F3\_C1530\_200). We run $32$ repetitions and use a budget of $200$ iterations for each task.
Table D.3: Hyperparameter spaces for ML models in TabRepo and TabRepoRaw.
ML model Hyperparameter Type Range Info Default value
NN(PyTorch) learning rate continuous [1e-4, 3e-2] log 3e-4
weight decay continuous [1e-12, 0.1] log 1e-6
dropout prob continuous [0, 0.4] 0.1
use batchnorm categorical False, True
num layers integer [1, 5] 2
hidden size integer [8, 256] 128
activation categorical relu, elu
NN(FastAI) learning rate continuous [5e-4, 1e-1] log 1e-2
layers categorical [200], [400], [200, 100],
[400, 200], [800, 400],
[200, 100, 50], [400, 200, 100]
emb drop continuous [0.0, 0.7] 0.1
ps continuous [0.0, 0.7] 0.1
bs categorical 256, 128, 512, 1024, 2048
epochs integer [20, 50] 30
CatBoost learning rate continuous [5e-3 ,0.1] log 0.05
depth integer [4, 8] 6
l2 leaf reg continuous [1, 5] 3
max ctr complexity integer [1, 5] 4
one hot max size categorical 2, 3, 5, 10
grow policy categorical SymmetricTree, Depthwise
LightGBM learning rate continuous [5e-3 ,0.1] log 0.05
feature fraction continuous [0.4, 1.0] 1.0
min data in leaf integer [2, 60] 20
num leaves integer [16, 255] 31
extra trees categorical False, True
XGBoost learning rate continuous [5e-3 ,0.1] log 0.1
max depth integer [4, 10] 6
min child weight continuous [0.5, 1.5] 1.0
colsample bytree continuous [0.5, 1.0] 1.0
enable categorical categorical False, True
Extra-trees max leaf nodes integer [5000, 50000]
min samples leaf categorical 1, 2, 3, 4, 5, 10, 20, 40, 80
max features categorical sqrt, log2, 0.5, 0.75, 1.0
Random-forest max leaf nodes integer [5000, 50000]
min samples leaf categorical 1, 2, 3, 4, 5, 10, 20, 40, 80
max features categorical sqrt, log2, 0.5, 0.75, 1.0
TabRepoRaw which uses the search space from TabRepo (details in Table D.3) and allows HPO to evaluate all configurations. For constructing TabRepo [Salinas and Erickson, 2024], each configuration was evaluated with a one-hour time limit and $8$-fold cross-validation. To reduce computational requirements for TabRepoRaw, we reduced this to $5$ minutes and $4$-fold cross-validation, and we provide it for $30$ datasets (context name: D244_F3_C1530_30D244\_F3\_C1530\_30). We use Bayesian optimization as implemented by SMAC [Lindauer et al., 2022] using Random Forests to conduct HPO. Additionally, we compare our two-level approach to combined search using SMAC, Random Search. We run $32$ repetitions and use a budget of $200$ iterations for the mentioned task.
To enable a fair comparison, we always evaluate the default configuration for each model first and then allow SMAC to run an initial design of 50#arms50-\#arms configurations in the upper level and $\frac{50}{\#arms-1}$ in the lower level.
Reshuffling [Nagler et al., 2024] which uses Heteroscedastic and Evolutionary Bayesian Optimization solver (HEBO) [Cowen-Rivers et al., 2022] for HPO. This benchmark includes HPO runs for $4$ ML models (details in Table D.4) across $10$ datasets, with $10$ repetitions and $3$ different validation split ratios within a budget of $250$ iterations. Although the benchmark does not support HPO over the entire search space, it offers a valuable opportunity to compare the performance of bandit methods in a realistic setting.
Table D.4: Hyperparameter spaces for ML models in Reshuffling.
ML model Hyperparameter Type Range Info
Funnel-Shaped MLP learning rate continuous [1e-4, 1e-1] log
num layers integer [1, 5]
max units categorical 64, 128, 256, 512
batch size categorical. 16, 32, …, max_batch_size
momentum continuous. [0.1, 0.99]
alpha continuous. [1e-6, 1e-1] log
Elastic Net C continuous [1e-6, 10e4] log
l1 ratio continuous [0.0, 1.0]
XGBoost max depth integer [2, 12] log
alpha continuous [1e-8, 1.5] log
lambda continuous [1e-8, 1.0] log
eta continuous [0.01, 0.3] log
CatBoost learning rate continuous [0.01 ,0.3] log
depth integer [2, 12]
l2 leaf reg continuous [0.5, 30]
D.3 Baselines and their hyperparameters
We use several bandit algorithms as baselines. Table D.5 summarizes the hyperparameters and their values.
Table D.5: Hyperparameters of Bandit Baselines.
Algorithm Hyperparameter Value Reference
MaxUCB $\alpha$ 0.5 Ours
Quantile Bayes UCB $\alpha$ 1.0 Balef et al. [2024]
$\beta$ 0.2
$\tau$ 0.95
Quantile UCB $\alpha$ 0.5
$\tau$ 0.95
ER-UCB-S $\beta$ 0.6 Hu et al. [2021]
$\theta$ 0.01
$\gamma$ 20.0
ER-UCB-N $\alpha$ 1.0
$\theta$ 0.01
$\gamma$ 20.0
Rising Bandits $C$ 7 Li et al. [2020]
$T$ Time horizon
Max-Median $\epsilon$ $1/(t)$, $t$ is iteration Bhatt et al. [2022]
QoMax-SDA $q$ 0.5 Baudry et al. [2022]
$\gamma$ 2/3
QoMax-ETC $q$ 0.5
$b_{T}$ 4
$n_{T}$ 3
$T$ Time horizon
UCB $\alpha$ 0.5 Auer [2002]
ThresholdAscent $\delta$ 0.1 Streeter and Smith [2006b]
$s$ 20
$T$ Time horizon
Successive Halving $\eta$ 2.0 Karnin et al. [2013]
$T$ Time horizon
R-SR $\epsilon$ 0.25 Mussi et al. [2024]
$T$ Time horizon
R-UCBE $\alpha$ 57.12
$\epsilon$ 0.25
$\sigma$ 0.05
$T$ Time horizon
MaxSearch (Gaussian) $c$ 1.0 Kikkawa and Ohno [2024]
MaxSearch (SubGaussian) $c$ 0.27
D.4 More Results on the Sensitivity Analysis of Hyperparameter $\alpha$
In addition to the results shown in Figure 5 in Section 5, we provide additional results here. We evaluated the performance of MaxUCB for $\alpha\in[0,2.9]$ with step size $0.1$. We plot the performance over the number of iterations for different values of $\alpha$ in Figures D.1, D.2, D.3, D.4. Green indicates better performance, showing the impact of $\alpha$ at different stages of the optimization procedure. Furthermore, we provide a comparison between MaxUCB with different values of $\alpha$ with combined search(SMAC) in Table D.6. For the experiments in the main paper, we choose $\alpha=0.5$ as a robust choice over all datasets. Notably, $\alpha=0.5$ is selected based on the assumption that the reward distribution support is $[0,1]$. For other supports, we recommend scaling $\alpha$ according to the range of the support.
Figure D.1: Heatmap showing the performance of our algorithm with different values of $\alpha$ for TabRepo dataset.
Figure D.2: Heatmap showing the performance of our algorithm with different values of $\alpha$ for TabRepoRaw dataset.
Figure D.3: Heatmap showing the performance of our algorithm with different values of $\alpha$ for YaHPOGym dataset.
Figure D.4: Heatmap showing the performance of our algorithm with different values of $\alpha$ for Reshuffling dataset.
Table D.6: Comparing MaxUCB with combined search(SMAC) for different values of $\alpha$ and time steps. P-values from a sign test assessing whether bandit methods outperform combined search. P-values below $\alpha=0.05$ are underlined, while those below $\alpha=0.05$ after multiple comparison correction (adjusting $\alpha$ by the number of comparisons) are boldfaced and underlined indicating that the two-level approach is superior to combined search. Additionally, we report the normalized loss and the number of wins, ties, and losses (w/t/l) of bandit methods.
Time $\alpha$=0.0 $\alpha$=0.1 $\alpha$=0.2 $\alpha$=0.3 $\alpha$=0.4 $\alpha$=0.5 $\alpha$=0.6 $\alpha$=0.7 $\alpha$=0.8 $\alpha$=0.9
TabRepoRaw[SMAC]
50 p-value $0.1002$ $\mathbf{0.0214}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$
w/t/l 1919/0/$11$ 2121/0/99 2727/0/33 2929/0/11 2929/0/11 2929/0/11 2929/0/11 2828/0/22 2727/0/33 2727/0/33
loss $0.2555$ $0.2446$ $0.2172$ $0.1946$ $0.1965$ $0.2134$ $0.2203$ $0.2229$ $0.2249$ $0.2296$
100 p-value $0.9974$ $0.9919$ $0.8998$ $0.1808$ $\mathbf{0.0081}$ $\mathbf{0.0214}$ $0.1002$ $0.5722$ $0.8192$ $0.9919$
w/t/l 88/0/$22$ 99/0/2121 1212/0/$18$ 1818/0/$12$ 2222/0/88 2121/0/99 1919/0/$11$ 1515/0/$15$ 1313/0/$17$ 99/0/2121
loss $0.2138$ $0.2053$ $0.1832$ $0.1346$ $0.1283$ $0.1113$ $0.1164$ $0.1208$ $0.1279$ $0.1512$
150 p-value $0.9993$ $0.9919$ $0.9506$ $0.2923$ $\mathbf{\underline{0.0007}}$ $\mathbf{\underline{0.0007}}$ $\mathbf{\underline{0.0026}}$ $\mathbf{0.0081}$ $0.1002$ $0.7077$
w/t/l 77/0/$23$ 99/0/2121 1111/0/1919 1717/0/$13$ 2424/0/66 2424/0/66 2323/0/77 2222/0/88 1919/0/$11$ 1414/0/$16$
loss $0.1994$ $0.1911$ $0.1671$ $0.1020$ $0.0898$ $0.0752$ $0.0775$ $0.0849$ $0.0887$ $0.0973$
200 p-value $0.9998$ $0.9993$ $0.9786$ $0.1808$ $\mathbf{0.0214}$ $\mathbf{\underline{0.0007}}$ $\mathbf{\underline{0.0026}}$ $\mathbf{\underline{0.0026}}$ $\mathbf{0.0081}$ $\mathbf{0.0081}$
w/t/l 66/0/$24$ 77/0/$23$ 1010/0/2020 1818/0/$12$ 2121/0/99 2424/0/66 2323/0/77 2323/0/77 2222/0/88 2222/0/88
loss $0.1864$ $0.1790$ $0.1489$ $0.0686$ $0.0651$ $0.0563$ $0.0622$ $0.0698$ $0.0703$ $0.0751$
YaHPOGym[SMAC]
50 p-value $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0000}}$
w/t/l 7474/11/$28$ 8383/11/1919 9595/11/77 102102/11/0 102102/11/0 102102/11/0 102102/11/0 102102/11/0 102102/11/0 101101/11/11
loss $0.1853$ $0.1532$ $0.1071$ $0.0930$ $0.0942$ $0.0978$ $0.1047$ $0.1101$ $0.1151$ $0.1190$
100 p-value $0.3833$ $0.3833$ $0.3833$ $\mathbf{0.0459}$ $\mathbf{0.0112}$ $\mathbf{0.0112}$ $\mathbf{0.0088}$ $0.0572$ $0.2153$ $0.4220$
w/t/l 5353/11/$49$ 5353/11/$49$ 5353/11/$49$ 6060/11/$42$ 6363/11/$39$ 6363/11/$39$ 6464/0/$39$ 6060/0/$43$ 5656/0/$47$ 5353/0/$50$
loss $0.1494$ $0.1177$ $0.0876$ $0.0813$ $0.0782$ $0.0713$ $0.0668$ $0.0702$ $0.0718$ $0.0749$
150 p-value $0.3833$ $0.6167$ $0.8135$ $0.0686$ $\mathbf{\underline{0.0036}}$ $\mathbf{0.0065}$ $\mathbf{\underline{0.0028}}$ $\mathbf{\underline{0.0005}}$ $\mathbf{\underline{0.0028}}$ $\mathbf{0.0148}$
w/t/l 5353/11/$49$ 5050/11/$52$ 4747/11/$55$ 5959/11/$43$ 6565/11/$37$ 6464/11/$38$ 6666/0/$37$ 6868/11/$34$ 6666/0/$37$ 6363/0/$40$
loss $0.1378$ $0.1013$ $0.0758$ $0.0722$ $0.0716$ $0.0551$ $0.0518$ $0.0507$ $0.0520$ $0.0500$
200 p-value $0.5394$ $0.8135$ $0.6896$ $0.2442$ $\mathbf{0.0185}$ $\mathbf{0.0088}$ $\mathbf{\underline{0.0015}}$ $\mathbf{\underline{0.0002}}$ $\mathbf{\underline{0.0000}}$ $\mathbf{\underline{0.0001}}$
w/t/l 5151/11/$51$ 4747/11/$55$ 4949/11/$53$ 5555/11/$47$ 6262/11/$40$ 6464/0/$39$ 6767/0/$36$ 7070/0/$33$ 7575/0/$28$ 7171/0/$32$
loss $0.1190$ $0.0856$ $0.0697$ $0.0660$ $0.0614$ $0.0457$ $0.0443$ $0.0418$ $0.0423$ $0.0421$
D.5 More Results for the Empirical Evaluation
In addition to the analysis in the main paper in Figure 6 in Section 5, we report the averaged normalized loss over time in Figure D.5, the average ranking in Figure D.6, the normalized loss per task in Figure D.7, the ranking per task in Figure D.8 and critical distance plots as described by Demšar [2006] in Figure D.9. Additionally, we report results for a Random Policy (yellow) that selects arms to pull at random.
TabRepo[RS] TabRepoRaw[SMAC] YaHPOGym[SMAC] Reshuffling[HEBO]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure D.5: Average normalized loss of algorithms on different benchmarks, lower is better. SMAC and random search perform combined search across the joint space.
TabRepo[RS] TabRepoRaw[SMAC] YaHPOGym[SMAC] Reshuffling[HEBO]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure D.6: Average ranking of algorithms on different benchmarks, lower is better. SMAC and random search perform combined search across the joint space.
Figure D.7: Heatmap showing the normalized loss of algorithms per task of each benchmark, sorted by the oracle arm performance, lower is better. SMAC and random search perform combined search across the joint space.
Figure D.8: Heatmap showing the ranking of algorithms per task of each benchmark, sorted by the oracle arm performance, lower is better. SMAC and random search perform combined search across the joint space.
TabRepo TabRepoRaw
Refer to caption Refer to caption
Refer to caption Refer to caption
YaHPOGym Reshuffling
Refer to caption Refer to caption
Figure D.9: Diagrams to compare the performance (ranking) of different algorithms using the Critical Distance (CD). For each benchmark on top, we compare bandit methods, and on the bottom, we compare MaxUCB against combined search and the oracle arm.
D.6 More Baselines for the Empirical Evaluation
Max $K$-armed Bandit Baselines. We compare MaxUCB against MaxSearch Gaussian [Kikkawa and Ohno, 2024], MaxSearch SubGaussian [Kikkawa and Ohno, 2024], QoMax-ETC [Baudry et al., 2022], QoMax-SDA [Baudry et al., 2022], Max-Median [Bhatt et al., 2022] and Threshold Ascent [Streeter and Smith, 2006b]. We report the averaged normalized loss over time in Figure D.10, the average ranking in Figure D.11. As shown, our algorithm outperforms all extreme bandit algorithms in these benchmarks.
TabRepo[RS] TabRepoRaw[SMAC] YaHPOGym[SMAC] Reshuffling[HEBO]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure D.10: Average normalized loss of MKB algorithms on different benchmarks, lower is better. SMAC and random search perform combined search across the joint space.
TabRepo[RS] TabRepoRaw[SMAC] YaHPOGym[SMAC] Reshuffling[HEBO]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure D.11: Average ranking of MKB algorithms on different benchmarks, lower is better. SMAC and random search perform combined search across the joint space.
A Few More Relevant Bandit Baselines. We compare MaxUCB against Quantile UCB [Balef et al., 2024], ER-UCB-N [Hu et al., 2021], R-SR [Mussi et al., 2024], R-UCBE [Mussi et al., 2024], Successive Halving [Karnin et al., 2013] and EXP3 [Auer et al., 2002]. We report the averaged normalized loss over time in Figure D.12, the average ranking in Figure D.13.
TabRepo[RS] TabRepoRaw[SMAC] YaHPOGym[SMAC] Reshuffling[HEBO]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure D.12: Average normalized loss of algorithms on different benchmarks, lower is better. SMAC and random search perform combined search across the joint space.
TabRepo[RS] TabRepoRaw[SMAC] YaHPOGym[SMAC] Reshuffling[HEBO]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure D.13: Average ranking of algorithms on different benchmarks, lower is better. SMAC and random search perform combined search across the joint space.
Appendix E More Details on the Empirical Behaviour of MaxUCB
Here, we provide further analysis of MaxUCB. Concretely, we study how often our algorithm pulls the optimal arm and compare it to the theoretical results. Furthermore, we evaluate an extension of MaxUCB to handle non-stationary rewards and finally study MaxUCB performance on common synthetic benchmarks used in the extreme bandit literature.
E.1 The number of times each arm is pulled
Proposition 4.1 shows that the number of times the optimal arm is pulled can be viewed as a good metric for measuring the performance of algorithms. Figure E.1, E.2,E.3,E.4 shows the average number of pulling arms on different benchmarks. They indicate that, on average, MaxUCB, Rising Bandits, and Max-Median algorithms often choose the optimal arm. However, for Max-Median, the number of pulls of the optimal arm is either very close to 0 or to $T$, leading to a non-robust performance, which has already been observed in Baudry et al. [2022] experiments. UCB and ER-UCB-S perform almost similarly.
Figure E.1: (Right) The number of all arm pulls, with each bar graph showing the average and the error bars indicating additional statistical information. (Left) The number of best arm pulls for different bandit algorithms.
Figure E.2: (Right) The number of all arm pulls, with each bar graph showing the average and the error bars indicating additional statistical information. (Left) The number of best arm pulls for different bandit algorithms.
Figure E.3: (Right) The number of all arm pulls, with each bar graph showing the average and the error bars indicating additional statistical information. (Left) The number of best arm pulls for different bandit algorithms.
Figure E.4: (Right) The number of all arm pulls, with each bar graph showing the average and the error bars indicating additional statistical information. (Left) The number of best arm pulls for different bandit algorithms.
E.2 From theory to practice
To validate our theorem against practical outcomes, we applied our algorithm to all benchmarks and plotted the number of pulls for each arm, denoted as "Real Experiment." Additionally, we computed the upper bound on the number of pulls by using the empirical values of $L_{1}$ and $U_{i}$ and $\Delta_{i}$. Notably, we report the first term of Equation 22 since the second term is nearly constant across all arms according to calculation. The results demonstrate that although the empirical pull counts are much less than the theoretical bounds of Equation 22, both follow a similar decreasing pattern as the rank of suboptimality increases.
Figure E.5: The number of pulls for each arm in our algorithm, labeled as "Real Experiment" and the theoretical values of this number
E.3 Addressing Non-stationary Rewards
To handle non-stationary rewards, we pull each arm $C$ times without observing the rewards before running MaxUCB. This "burn-in" allows the Markov Chain to reach equilibrium, especially from a poor starting point. Algorithm E.1 shows the adapted version of our algorithm. Therefore, empirically, this allows MaxUCB to operate after a fixed exploration phase of all arms until the reward distribution is stationary.
We run Algorithm E.1 with different parameters of $C\in\{5,6,7,8\}$ using up to $48$ iterations corresponding to almost $25\%$ of the total budget. Figure E.6 shows normalized loss per task where columns are sorted by the maximum change of the mean of the reward distributions of the optimal arm computed every $10$ HPO iterations (as an indicator of non-stationarity; shown at the top panel in Figure E.6. Figure E.7 shows the average ranking and normalized loss over time for different values of the hyperparameter $C$.
The initial burn-in improves final performance for the few tasks where we observe a high shift (right part of Figure E.6) while the initial performance is worse across all tasks (as shown in Figure E.7). However, the results are not sensitive to the exact value of $C$. Overall, this naive solution can improve performance for some tasks at the cost of not using potentially valuable information obtained from initial exploration. Thus, optimally addressing non-stationary rewards could be a promising direction for future work.
Algorithm E.1 MaxUCB-Burn-in
[…]
2:for $j\leq C$, for each arm $i\leq K$ do
3:  Pull arm $i$
[…]
5:for each arm $i\leq K$ do
6:  Pull arm $i$
7:  set $n_{i}\leftarrow 1$, observe reward $r_{i,1}$
[…]
9:for $t=(CK+K+1)$ to $T$ do
10:  for each arm $i\leq K$ do
11:    Update policy $U_{i}=\max{(r_{i,1},...,r_{i,n_{i}})}+(\frac{\alpha\log(t)}{n_{i}})^{2}$
12:  end for
13:  Select arm $I_{t}=\operatorname*{arg\,max}\limits_{i\leq K}U_{i}$
14:  $n_{I_{t}}\leftarrow n_{I_{t}}+1$
15:  Observe reward $r_{I_{t},n_{I_{t}}}$
16:end for
Figure E.6: Heat map shows normalized loss per task, sorted by the distribution shift.
Figure E.7: Average rank and normalized loss of algorithms on YaHPOGym benchmark, lower is better.
E.4 Toy examples from the extreme bandit’s literature
In this section, we provide additional results on commonly used benchmark functions used in the extreme bandit literature. Concretely, we use a similar setup to [Baudry et al., 2022] and report the following four tasks:
  1. 1.
    $K=5$ Pareto distributions with tail parameters parameters $\lambda_{k}=[2.1,2.3,1.3,1.1,1.9]$. Results are shown in Figure E.8.
  2. 2.
    $K=10$ Exponential arms with a survival function $G_{k}(x)=e^{-\lambda_{k}x}$ with parameters $\lambda_{k}=[2.1,2.4,1.9,1.3,1.1,2.9,1.5,2.2,2.6,1.4]$. Results are shown in Figure E.9.
  3. 3.
    $K=20$ Gaussian arms, with same mean $\mu_{k}=1,\forall k$, and different variances $\sigma_{k}=[1.64,2.29,1.79,2.67,1.70,1.36,1.90,2.19,0.80,0.12,1.65,1.19,1.88,0.89,3.35,1.5,2.22,3.03,1.08,0.48]$. The dominant arm has a standard deviation of $3.35$. Results are shown in Figure E.10.
  4. 4.
    (our toy example) $K=5$ power distributions with domain parameter $[3,4,5,5,4]$ and shape parameter $[1.01,1.01,1.01,1.1,1]$. Results are shown in Figure E.11.
For each task, we run $N=1000$ independent repetitions for six time horizons $T\in\{50100,200,500,1000,2000\}$. We show the CDF of the rewards for each arm, the number of times the optimal arm was pulled, and the proxy empirical regret. Notably, the proxy empirical regret is introduced by Baudry et al. [2022] to overcome the issue of high variance in the maximum values of distributions.
Figure E.8: Experiment 1: (Top) Survival function of distribution of each arm (left) Number of pulls of the optimal arm. (Right) Proxy Empirical Regret
Figure E.9: Experiment 2: (Top) Survival function of distribution of each arm (left) Number of pulls of the optimal arm. (Right) Proxy Empirical Regret
Figure E.10: Experiment 3: (Top) Survival function of distribution of some arms (left) Number of pulls of the optimal arm. (Right) Proxy Empirical Regret
Figure E.11: Experiment 4: (Top) Survival function of distribution of each arm (left) Number of pulls of the optimal arm. (Right) Proxy Empirical Regret
E.5 Supernet selection in Few-Shot Neural Architecture Search
In one-shot NAS, a single supernet approximates all architectures. However, this estimation could be inaccurate. To address this, few-shot NAS splits the supernet into smaller sub-supernets [Hu et al., 2022, Ly-Manson et al., 2024]. To further improve performance, [Hu et al., 2022] introduced supernet selection, which identifies the most promising sub-supernet and its optimal architecture using techniques such as Successive Halving . We aim to identify the best-performing architecture among several subspaces, each corresponding to a sub-supernet. This problem is analogous to the MKB problems, where each arm represents a sub-supernet.
Dataset Preparation. We use data provided in [Ly-Manson et al., 2024] for three benchmark datasets: CIFAR-10, CIFAR-100, and ImageNet16-120. Each dataset’s search space is split using $10$ different metrics. The splitting follows a binary tree structure with a depth of $3$, where operations are divided into two groups at each branch. This process results in $8$ sub-supernets per metric. Consequently, for each dataset, we have one full search space and $8$ sub-search spaces.
Each combination of dataset and splitting metric is treated as a separate task, yielding a total of $30$ tasks ($3$ datasets × $10$ metrics), each with $8$ arms. Following the setup of [Ly-Manson et al., 2024], we randomly sample $600$ architectures from both the full search space and each sub-search space. This process is repeated $32$ times using different random seeds to ensure variability and robustness in the results.
Analyzing the reward distribution. Figure E.12 illustrates the empirical survival functions of the rewards and sub-optimality gaps for the benchmark. As shown, the distribution shape is similar to that of HPO tasks: both are bounded and left-skewed. However, the sub-optimality gap is considerably smaller than HPO tasks, suggesting that identifying the optimal arm is more challenging and may require additional iterations. In Figure E.13, we show values of $L$ and $U$ from Lemma 3.3 for this benchmark.
Supernet
Refer to caption Refer to caption Refer to caption
Figure E.12: (Left) The average empirical survival function of the reward (observed performances) per arm ranked per dataset. We divided the reward sequence into five segments over the budget to show the distribution change over time. Thin lines correspond to the survival function of different segments, visualizing the change over time. (Middle) The average ECDF per dataset for the best and worst arm with thin lines corresponding to individual datasets. (Right) The sub-optimality gap $\Delta_{i}$.
Supernet
Figure E.13: Arms are ordered by sub-optimality gap. (Top) Thin orange lines represent $\frac{G(b-\epsilon)}{\epsilon}$, while the blue and red points correspond to $L$ and $U$ for our empirical reward distributions (see Lemma 3.3 for details). (Middle) Histogram of values for $L$. (Bottom) Histogram of values for $U$.
Performance Analysis. Figure E.14 presents the average ranking and normalized loss of various bandit algorithms in this benchmark. As shown, Successive Halving, Max-Median, and ER-UCB-S perform well with a small time budget but fail to explore sufficiently to identify the optimal arm. Rising Bandits, as a fixed-confidence best-arm identification method, struggles to find the optimal arm. In contrast, MaxUCB outperforms all other baselines, demonstrating better performance when searching the entire search space with a higher budget.
Figure E.14: Average rank and normalized loss of algorithms on Supernet Selection benchmark, lower is better.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We wrote the abstract and introduction to reflect the paper’s contribution (a practical Bandit strategy for decomposed CASH) as detailed as possible and discuss it ability to generalize to other settings in the Conclusion (Section 6) and in Appendix B.4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide a separate paragraph for the limitations of our methods in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Yes, we provide all necessary assumptions for the theoretical results in Section 3 and detail proofs for theories in Appendix B. We also discuss the validity of our assumptions using on empirical data analysis in Section 3.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, we provide detailed information about all the methods we used. This includes a description of the computing infrastructure in Footnote 5, an ablation study for hyperparameter selection in Section 5 and Appendix D.4, as well as metrics and experimental details in Appendix D.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.checklist
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We make our code and data available at https://anonymous.4open.science/r/CASH_with_Bandits/README.md.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, in addition to making code available, we provide detailed information about all the methods we used, including the choice of hyperparameters. This includes a description of the computing infrastructure in Footnote 5, an ablation study for hyperparameter selection in Section 5 and Appendix D.4, as well as metrics and experimental details in Section D.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report confidence intervals concerning the random seed after running the experiments multiple times (see Figure 6 and Table 1 in the main paper), and we discuss this further in the Appendix D.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide a description of the computing infrastructure in Footnote 5.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our work complies with the NeurIPS Code of Ethics. It involves no human subjects, sensitive data, or misuse risks, and uses only public datasets.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: This paper focuses on advancing the field of Machine Learning without direct societal impacts or specific applications that would need to be discussed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work poses no known risks of misuse and does not involve high-risk models or datasets.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All external assets (datasets and code) used in our work are publicly available, properly cited, and used in accordance with their respective licenses.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented, and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not introduce any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This work does not involve crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This research does not involve human subjects or require IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We only used LLMs for text polishing and coding assistance.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.